Three masses are located in the corners of an equilateral triangle. Find the magnitude and direction of the gravitational field at the center of the triangle. Given: m₁ = 22 kg, m₂ = 30 kg, m₃ = 30 kg, r = 12 cm. G = 6.674×10–11 N·m²/kg². 


1
Expert's answer
2021-03-21T11:25:59-0400

Let "r" is the distance between the masses and the center of an equilateral triangle. So, we have


"E_1=G\frac{m_1}{r^2}=6.674\cdot10^{-11}\cdot \frac{22}{0.12^2}\approx1.02\cdot10^{-7}\ (N/kg)"


"E_2=G\frac{m_2}{r^2}=6.674\cdot10^{-11}\cdot \frac{30}{0.12^2}\approx1.39\cdot10^{-7}\ (N/kg)"


"E_3=G\frac{m_3}{r^2}=6.674\cdot10^{-11}\cdot \frac{30}{0.12^2}\approx1.39\cdot10^{-7}\ (N/kg)"


"|\vec E_2+\vec E_3|=2\cdot1.39\cdot10^{-7}\cdot\cos60°=1.39\cdot10^{-7}\ (N/kg)"


The gravitational field at the center of the triangle


"E=1.39\cdot10^{-7}-1.02\cdot10^{-7}=0.37\cdot10^{-7}\ (N/kg)" . Answer